OpenStudy (anonymous):

someone HELP ME: write the expression as a complex number in standard form. (-1+5i)(6+3i)

OpenStudy (zzr0ck3r):

-1*6+(-3i)+30i+15&i^2 -6+27i-15 27i-21

OpenStudy (mathstudent55):

Here you are multiplying complex numbers. You start out by multiplying with FOIL. (-1 + 5i)(6 + 3i) = = -1 * 6 + (-1) * (3i) + 5i * 6 + 5i * (3i) Now you do the multiplications: = -6 - 3i + 30i + 15i^2 Now you simplify i^2, since i^2 = -1 = -6 - 3i + 30i + 15(-1) Multiply 15 by -1 = -6 - 3i + 30i - 15 Now you combine like terms like we did before. = -6 - 15 - 3i + 30i = -21 + 27i This is standard form for a complex number

OpenStudy (zzr0ck3r):

same thing as(a+b)(c+d) but i^2=-1

OpenStudy (anonymous):

 write the expression as a complex number dividing complex numbers